4n + 3 = 5(2 – 2n)
Answer by kkasko(46)

4n + 3 = 5(2 – 2n)
4n+3=10-10n
move -10 to the left side and move 3 to the right side by doing the opposite
4n+3=10-10n
+10n -3. -3.
14n=7
/14. /14
n=7/14
Reduce
7/14=1/2
n=1/2